A company must pay liabilities of 3000 and 5000 at the end of years 2 and 4, respectively. The only investments available to the company are thre following two zero - coupons bonds : 1) maturity 2 years, effective annual yield 5.5%, par 1000. 2) maturity 4 years, effective annual yield 6.8%, par 100. Find the cost of exactly matching those liabilities.
Answer #1

The company must invest the present values of 3000 in 2 years at 5.5% and 5000 in 4 years at 6.8%.

The cost is given by= 3000/(1.055^2) + 5000/(1.068^2) = 2695.36 + 4383.57 = 7078.93